OpenStudy (faiithhh):

solve the system by substitution {-x-y-z=-8 -4x+4y+5z=7 2x+2z=4}

OpenStudy (jango_in_dtown):

2x+2z=4 so (x+z)=2

OpenStudy (nevillesnetwork):

is this a multiple choice question, are there answers to choose from?

OpenStudy (jango_in_dtown):

put this in the first equation -y-(x+z)=-8i.e. -y-2=-8 i.e. y=8-2=6

OpenStudy (nevillesnetwork):

jango is correct

OpenStudy (faiithhh):

no, not multiple choice. i have to show my work, but i don't know how to do any of this.

OpenStudy (jango_in_dtown):

then use the results in equation 2 , which you can write -4x+4*6+5z=7i.e. -4x+5z=7-24=-17 i.e. 4x-5z=17

OpenStudy (jango_in_dtown):

now we need to solve x+z=2 and 4x-5z=17,

OpenStudy (faiithhh):

okay... woah... i mean like, i literally don't know how to do any of this. Everything you've said has gone right over my head...

OpenStudy (jango_in_dtown):

Do you know what is the method of substitution?

OpenStudy (faiithhh):

not really, no..

OpenStudy (jango_in_dtown):

Then you must first learn the theory. :)

OpenStudy (faiithhh):

could you teach me, oh great & wise one? xD